# Gardeyn, Vanden Berghe and Wauters

### v1

@@ 1 Introduction
While the objectives and particulars of these irregular C&P nesting problems can differ, they all share the need for a specific feasibility check: determining whether or not an item can be cut from or placed at a certain position without overlapping with anything else. Unlike, say, the traveling salesman problem in which feasibility simply means ensuring all locations are visited, here the irregularity of the items and containers makes implementing this check in a precise and efficient manner a challenging endeavor. Therefore, in addition to the optimization challenge of finding a good solution according to some objective function, irregular C&P problems also pose an additional geometric challenge of accurately and efficiently verifying the feasibility of candidate solutions.
@@ 2 General Approaches to Geometry in 2D C&P Problems
The most significant challenge for this approach undoubtably remains the development of robust NFP generators, particularly for non-convex and complex shapes. Over the years, numerous papers have been published claiming to improve NFP generation robustness or efficiency. Yet, at the time of writing this article, a publicly available implementation without significant shortcomings remains elusive. Even the authors of the most popular NFP generator on GitHub (Mészáros (2018-2022)) acknowledge that their implementation is not feature-complete and only works well for shapes without holes or concavities.
@@ 3 Collision Detection Engine
First, we can assume a dense environment (unlike video games), which means that the probability of collision is high. Second, we can assume a largely static environment (unlike simulations), which means that queries will be more frequent than updates. Third, we are only interested in the boolean answer to the query. Therefore we only need to determine whether a collision occurs, but not necessarily the details concerning where and how exactly it occurs.
@@ 4 Polygon Collision
Figure 1(b) illustrates the distinct types of collision that exist between two simple polygons. The first type of collision occurs when the polygons have edges that intersect each other. If there are no intersecting edges, the two polygons are either disjoint or one polygon is located entirely inside the bounds of the other. The second type of collision accounts for the second case, where one polygon is fully contained within the other. Every collision between two simple polygons can be expressed as one of these two types (Preparata and Shamos (1985)).
@@ 7.1 Poles (of inaccessibility)
Agafonkin (2016) proposed an efficient algorithm to calculate a polygon’s PoI up to a certain (configurable) precision. For use in the proposed surrogate, we extended Agafonkin’s algorithm so that it can iteratively generate multiple non-overlapping poles. Each iteration generates the next pole, with all poles from previous iterations being treated as holes in the shape. Figure 5(a) shows the first five poles the algorithm generates for the bunny.
@@ 7.2 Piers
The main body of the bunny in Figure 5(a) is captured quite well by the five poles, but its ears and legs are not represented at all. Long and narrow sections are particularly challenging to capture with a limited set of circles. This is not ideal since these “extremities” are quite likely to collide with hazards. To combat this, we include a second component of surrogate representation to complement the poles.
@@ 9 Codebase
The CDE implementation is called jagua-rs and its source code is available at: https://github.com/JeroenGar/jagua-rs. We wrote this project entirely in Rust: a statically-typed and compiled programming language with a strong emphasis on performance, safety and robustness. We have currently implemented both strip- and bin-packing problems, however the project can be extended to support virtually any 2D irregular C&P problem variant.
@@ 10 Validation & evaluation
In our opinion, the only irrefutable way to validate (i) the use of trigonometry, (ii) the general design of the CDE and (iii) its implementation, is to build a competitive optimization algorithm on top of jagua-rs. Readers interested in seeing jagua-rs in action are referred to the sparrow project: an open-source heuristic for the 2D strip packing problem (Gardeyn (2025)). Built on top of jagua-rs, sparrow demonstrates the CDE’s capabilities in practice.
For readers interested in a relative performance evaluation, we use the criterion.rs framework by Heisler and Aparicio (2014–2024) to perform “statistics-driven microbenchmarking” of the most critical operations within the CDE. We encourage readers interested in using the CDE to re-run the tests on their own hardware and problem instances using the latest version of jagua-rs. These benchmarks should provide valuable insights into the factors at play and the trade-offs involved when tuning the CDE for maximum performance.
@@ 11 Conclusion
Our paper introduced a suite of core principles and design philosophies which we believe are essential when modelling a CDE for use in irregular C&P problems: hazards, two-phased approach, fail-fast surrogate and polygon simplification. However, such concepts are only useful insofar as they can actually be leveraged to achieve an effective implementation.

### v4

@@ 1 Introduction
While the objectives and particulars of these irregular C&P — nesting — problems can differ, they all share the need for a specific feasibility check: determining whether or not an item can be cut from or placed at a certain position without overlapping with anything else. Unlike, say, the traveling salesman problem in which feasibility simply means ensuring all locations are visited, here the irregularity of the items and containers makes implementing this check in a precise and efficient manner a challenging endeavor. Therefore, in addition to the optimization challenge of finding a good solution according to some objective function, irregular C&P problems also pose an additional geometric challenge of accurately and efficiently verifying the feasibility of candidate solutions.
@@ 2 General Approaches to Geometry in 2D C&P Problems
The most significant challenge for this approach undoubtably remains the development of robust NFP generators, particularly for non-convex and complex shapes. Over the years, numerous papers have been published claiming to improve NFP generation robustness or efficiency. Yet, at the time of writing this article, a publicly available implementation without significant shortcomings remains elusive. Even the authors of the most popular NFP generator on GitHub (Mészáros (2018)) acknowledge that their implementation is not feature-complete and only works well for shapes without holes or concavities.
@@ 3 Collision Detection Engine
First, we can assume a dense environment (unlike video games), which means that the probability of collision is high. Second, we can assume a largely static environment (unlike simulations), which means that queries will be more frequent than updates. Third, we are only interested in the boolean answer to the query. Therefore, we only need to determine whether a collision occurs, but not necessarily the details concerning where and how exactly it occurs.
@@ 4 Polygon Collision
Figure 1(b) illustrates the distinct types of collision that exist between two simple polygons. The first type of collision occurs when the polygons have edges that intersect each other. If there are no intersecting edges, the two polygons are either disjoint or one polygon is located entirely inside the bounds of the other. The second type of collision accounts for the second case, where one polygon is fully contained within the other. Every collision between two simple polygons can be expressed as one of these two types Preparata and Shamos (1985).
@@ 7.1 Poles (of inaccessibility)
Agafonkin (2016) proposed an efficient algorithm to calculate a polygon’s PoI up to a certain (configurable) precision. For use in the proposed surrogate, we extended their algorithm so that it can iteratively generate multiple non-overlapping poles. Each iteration generates the next pole, with all poles from previous iterations being treated as holes in the shape. Figure 5(a) shows the first five poles the algorithm generates for the bunny.
@@ 7.2 Piers
The main body of the bunny in Figure 5(a) is captured quite well by the five poles, but its ears and legs are not represented at all. Long and narrow sections are particularly challenging to capture with a limited set of circles. This is not ideal since these ‘extremities’ are quite likely to collide with hazards. To combat this, we include a second component of surrogate representation to complement the poles.
@@ 9 Codebase
The CDE implementation is called jagua-rs and its source code is available at: https://github.com/JeroenGar/jagua-rs, with a snapshot archived in the INFORMS Journal on Computing repository (Gardeyn 2025). We wrote this project entirely in Rust: a statically-typed and compiled programming language with a strong emphasis on performance, safety and robustness. We have currently implemented both strip- and bin-packing problems, however the project can be extended to support virtually any 2D irregular C&P problem variant.
@@ 10 Validation & evaluation
In our opinion, the only irrefutable way to validate (i) the use of trigonometry, (ii) the general design of the CDE and (iii) its implementation, is to build a competitive optimization algorithm on top of jagua-rs. Readers interested in seeing jagua-rs in action are referred to the sparrow project: an open-source heuristic for the 2D strip packing problem (Gardeyn et al. (2025)). Built on top of jagua-rs, sparrow demonstrates the CDE’s capabilities in practice.
For readers interested in a relative performance evaluation, we use the criterion.rs framework by Heisler and Aparicio (2014) to perform “statistics-driven microbenchmarking” of the most critical operations within the CDE. We encourage readers interested in using the CDE to re-run the tests on their own hardware and problem instances using the latest version of jagua-rs. These benchmarks should provide valuable insights into the factors at play and the trade-offs involved when tuning the CDE for maximum performance.
@@ 11 Conclusion
Our paper introduced a suite of core principles and design philosophies which we believe are essential when modeling a CDE for use in irregular C&P problems: hazards, two-phased approach, fail-fast surrogate and polygon simplification. However, such concepts are only useful insofar as they can actually be leveraged to achieve an effective implementation.